Session: Feb 1999 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Mantovani at Wed Feb 17 09:17:46 MET 1999) Success
No known problem

Torun (K.Borkowski at Wed Feb 17 10:00:21 MET 1999) Success
First 8 minutes of scan at 18:56 lost due to antenna having to rotate through more than 300 deg in the azimuth. It was snowing (intermitantly) during the experiment.

Onsala (A. Polatidis at Wed Feb 17 10:10:45 MET 1999) Success
Snow during the first few hours of the experiment. Cloudy during the remaining part. No obvious problems with the data.

Westerbork (Tony Foley at Wed Feb 17 14:08:46 MET 1999) Failure
from start (18:00) to 22:00 looks OK, but very heavy snowstorm Extinction about 100% from about 22:00, dropping b04ack to noraml at 04:00UT. Telescope drive frozen with large lump of ice at about 02:00UT (only removed during daylight)

Jodrell (Alastair Gunn at Fri Feb 19 14:13:43 MET 1999) Minor failures
Telescope parked due to high winds from start of experiment until 02:19 UT. Only tape 3 sent to correlator. No other known problems.

Cambridge (Alastair Gunn at Fri Feb 19 14:15:37 MET 1999) Minor failures
The procedure file specifies IFs c,a,c,a but we normally run as a,c,a,c. May need to flip the polarisations. The source change at 14:03 UT did not succeed due to MERLIN problems. No other known problems.

Effelsberg (dgraham at Mon Feb 22 15:26:37 MET 1999) Minor failures
stopped due to snow 0410-0711. Azm stopped for a short time around 2237.

Noto (C.Trigilio at Mon Feb 22 23:18:55 MET 1999) Success

Calibration (Jean\-Francois Desmurs at Wed Feb 24 11:55:14 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: ew010.antab. Read the comments in the file to get further information.